May 29, 1996

Atomic Databases

Name  Country  Contact  Access  Data
TOPbase  France  C.Mendoza  Telnet  EL,OS,TP,PH,WL
NIFS  Japan H.Tawara,T.Kato  Telnet  EX,IN,CT,HP
NIST ASD  U.S.A.  W.C.Martin  WWW,Telnet EL,OS,TP,WL
CfA Harvard  U.S.A.  P.L.Smith  WWW  EL,OS,TP,WL
CDS  France  —  FTP  VSC
CFADC  U.S.A.  D.R.Schultz  WWW  IN,Bibl: VSC
GAPHYOR  France  J.L.Delcroix  WWW,E-mail  Bibl: VSC
IAEA AMDIS Austria  R.K.Janev  Telnet  EX,IN,CT,HP,Bibl
VALD  Austria  F.Kupka  E-mail  EL,OS,TP
CCP7  U.K.  C.S.Jeffery  FTP  EL,WL
SAM  U.S.A.  T.Brage  WWW  EL,OS,TP,HS
DASGAL  France  A.Lesage  WWW  Bibl: LS,SH
ADA  U.S.A.  D.Verner  WWW  VSC
NIST ATPBD U.S.A.  P.J.Mohr  WWW  Bibl: OS,TP
ADRAL  Canada  D.C.Morton  WWW  EL,OS,TP
BIBL  Russia  A.N.Ryabtsev  FTP  Bibl: VSC
LANL EPAS  U.S.A.  L.Kissel  WWW  SF
Uppsala U.  Sweden  —  WWW  SF,WL,OS
ADB Lund U. Sweden  A.Contis  WWW  EL,WL,OS
RTAM  Finland  P.Pyykko  WWW  Bibl: VSC

Data keys:

Bibl: XXX - bibliography for XXX; CT - charge transfer cross sections; EL - energy levels; EX - electron impact excitation cross sections; HP - heavy-particles interaction cross sections; HS - hyperfine structure parameters; IN - electron impact ionization cross sections; LS - line shapes; OS - oscillator strengths; PH - photoionization cross sections; SH - line shifts; SF - scattering factors; TP - transition probabilities; VSC - various spectroscopic and collisional data; WL - wavelengths.